OpenStudy (anonymous):

studying the function derivable f: \[(0, + \infty)\] ....>R \[\sqrt{3+x ^{2}} - x if 0 < x \le 1 \] \[\frac{ 1 }{ x } +\frac{ x ^{2} }{4} if x > 1\]

OpenStudy (anonymous):

@hartnn @hba

hartnn (hartnn):

to check whether the function is differentiable at x=1, check whether its continuous at x=1. if its not continuous, then its not differentiable at x=1

hartnn (hartnn):

know how to find whether its continuous or not ?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} 3 +x ^{2} - x = 3+1-1= 3\] \[\lim_{x \rightarrow 1} 1/x + x ^{2}/4= 1/1 + 1/4 = 5\]

OpenStudy (anonymous):

is not continuous

hartnn (hartnn):

so, its not differentiable at x=1

OpenStudy (anonymous):

is it correct?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

 how can solve